# h FR-PHENO-2025-02 A general approach to quantum integration of cross sections in high-energy physics

### v1

@@ 2.4 Implementing a generic cross-section calculation
In the introduction, we have mentioned so called generic building blocks for the computation of cross sections. To illustrate this, we provide here two concrete examples. From Eq. 1 and Eq. 4, a cross section with only a single integration variable reduces to
@@
where $s$ is the maximal value of $x$. In order to cast the above equation into the more rigorous one of the MCI problem as in Eq. 6, noting that we are free to choose how to split the integrand for the expectation value into the product of function applied, $g(x)$, and probability distribution, $f_{X}(x)$, it follows that we can choose the function applied to be
@@
(i.e., the numerator). The probability distribution to sample from can then be associated to the propagator function (i.e., the denominator), such that
@@
The reason we chose this particular factorisation is twofold. First, as mentioned previously, implementing moments or products of moments of random variables as the function applied is straightforward to implement based on the QMCI engine’s existing functionality, and the RMSE scaling of the corresponding Fourier series’ are likely favourable as compared to those of more complex functions. Second, because the probability distribution corresponds to a real distribution that has physical significance (being used to model resonance behaviour in general), we can envision that future development of, for example, specific state-preparation methods tailored to this distribution [^8] Note that in this article we only consider general state-preparation methods for preparing the BW distribution, and we defer research into more tailored methods for future work. would not only be useful for this particular application, but also for future research directions in HEP (e.g., for quantum-computing applications where resonance behaviours are required to be modelled).
We can identify the general form of Eq. 14 as an (unnormalised) relativistic Breit-Wigner (BW) distribution [46] (hereafter simply denoted by BW distribution). As mentioned previously, state preparation of arbitrary probability distributions is a difficult problem. However, in this case we need only prepare one specific distribution, and as will be discussed in the following section, it will suffice to prepare this distribution a single time (for a given COM energy) for each of the (small number of) relevant massive particles in the Standard Model, i.e., the W and Z bosons, the top (t) quark, and the Higgs boson. In the present example, the two building blocks are the function, $g$, which is a monomial, and the BW distribution, $f_{X}$.
@@
Such terms would then be absorbed in the monomials in the numerator, possibly with negative exponents. For this case of negative exponents, because the functions applied for FQMCI must obey certain smoothness conditions, as detailed previously in Section 2.3.1, then in this case one would not be able to exploit FQMCI. Instead, an approach similar to those of prior works would be required, where the entire integrand is prepared as the probability distribution and then QAE run directly on this state (equivalent in our language to the “function applied” being constant).
@@
where the term $x^{n}y^{m}$ means that the numerator can be made of any monomials which are a product of $x$ and $y$ raised to various powers. It implies that in order to match the form of Eq. 8, the separable functions applied should be
@@
The QMCI engine provides efficient Fourier decomposition circuits for moments or products of moment of random variables, as discussed previously, and allows the user to provide the probability distribution to be sampled from as an input. In addition, when constraints on the phase space are required to be imposed on the integration, as given by the cut function, $C$, in Eq. 5, then for simple constraints such as sums, maximum/minimum etc., the engine’s built-in thresholding operations can be used to achieve this. Thus—providing there are efficient ways to prepare BW distributions as quantum states, something discussed in the next section—in principle the QMCI engine can be used to perform quantum integration of a generic cross section in HEP.
It is worth noting that the cut function, $C$, is a function of the integration variables, and as discussed previously, possibly non-analytical. This means that in order to implement the restriction of phase space, the value $C(\Phi)$ should be numerically evaluated at each phase-space point. Such an approach has been followed in Section 4 to set the boundaries of integration of a given variable that depends on another variable of integration; however, in this case the expression was analytical, and the quantum arithmetic implemented via the QMCI engine’s enhanced $P$-builder. For the case of some non-analytical form, one possibility is to implement the required quantum arithmetic by constructing the equivalent classical circuit, and then making it a reversible circuit. However, we note that this approach is unlikely to be efficient, and indeed, in the future there may be better ways of constructing such circuits. We leave this aspect for future work.
3 State preparation of relativistic Breit-Wigner distributions
We see that to be able to perform generic cross-section calculations using the QMCI engine, we require a methodology for preparing on a quantum register BW distributions for the massive propagators W, Z and t, [^9] In practice, completeness would also require a propagator for the Higgs boson. However, due to the narrow width of the resonance, generating this peak with sufficient resolution is highly challenging, given current qubit counts and methods. At high-energy colliders, all leptons are typically considered as massless, alongside all quarks (apart from the top quark), as their masses are negligible at these energy scales. In certain cases, the mass of the bottom quark is also included in theoretical calculations, but with the width assumed to be zero. as these constitute one of the building blocks of the method. As discussed briefly in the introduction, state preparation of probability distributions for QMCI is an active research topic, and several methods have been proposed in the literature (see, e.g., Refs. [47, 48, 49, 50, 51, 52, 53]); however, there is currently no standard methodology that is commonly used.
@@ 3 State preparation of relativistic Breit-Wigner distributions
Before we continue the discussion, however, we must first discuss some of the challenges inherent to state preparation for QMCI. These challenges generally arise due to the limited number of qubits, $n$, that are available (due to the limitations of current hardware) to represent the true continuous distribution as a discretised, truncated distribution. In the QMCI engine, the amplitudes of the state, |$\ket{p}$, are interpreted as the support points of the discretised distribution, with the number of support points, $N=2^{n}$. These are uniformly spaced with spacing size, $\Delta$, spanning the range of the support of the distribution, $[x_{l},x_{u}]$. In practice, the limited number of points alongside the uniform binning leads to limited resolution on the distribution of interest, as well as having to truncate what could be an infinite distribution. For a finite number of qubits, these limitations introduce a number of systematic errors into a QMCI calculation. [^10] It is important to note that many of these systematic errors are also applicable to classical MCI, as will be touched upon. Reference [33] discusses these systematic errors in detail, and so here we merely reproduce some of the definitions of the relevant ones for this discussion:
@@
  • State-preparation error: this error is distinct from the other errors discussed as it does not arise from the limited number of qubits, but rather the imperfect method of preparing the distribution as a quantum state i.e., there is a difference between the ideal quantum state and the approximated state that is actually prepared. The state-preparation error can be defined in a variety of different ways. In this article, we consider the cumulative-distribution-function (CDF) mean-squared error (CMSE) between the probabilities of the prepared state, $f^{p}_{X}(x_{i})$, and the true probabilities, $\tilde{f}_{X}(x_{i})$, as the metric representing the state-preparation error
@@
All of these systematic errors can be considered relevant for the cross-section calculations considered later in Section 4.
Considering the discretisation and normalisation errors in particular, for a given number of qubits, $n$, there will be a sub-range of the full potential support of the distribution where these systematic errors are sufficiently minimised (below a chosen bound) to have a negligible impact on a QMCI calculation. When considering cross-section calculations for a given experiment, the integration is only ever performed from $s_{min}=0\,\text{GeV}^{2}$ to some $s_{max}=S\,\text{GeV}^{2}$. Therefore, in practice, it will not be necessary to prepare the BW distribution to span the full potential support range [$s_{min}=0\,\text{GeV}^{2}$ to $s_{max}=+\infty\,\text{GeV}^{2}$]—which would require an infinite number of qubits in order to sufficiently suppress these errors. Thus, whilst qubit numbers are still limited, for practical applications it will instead suffice to generate a range of circuits that prepare BW distributions with supports spanning a range of different COM energies squared, with each prepared using a sufficient number of qubits as to sufficiently suppress these systematic errors.
[…]
To this end, and for illustration purposes, in this article we consider preparing BW distributions for the W, Z, and t resonances, corresponding to a COM energy, $\sqrt{s_{max}}=200\,\text{GeV}$, and for just the W to $\sqrt{s_{max}}=100\,\text{GeV}$. The value, $\sqrt{s_{max}}=100\,\text{GeV}$, was chosen such that the circuits for the cross-section calculations in Sections 4.2 and 4.3 (which only use the W-boson propagator) were sufficiently small to be classically simulable using state-vector methods, while $\sqrt{s_{max}}=200\,\text{GeV}$ was chosen in order to be able to generate distributions covering all of the resonances.
Figures 1 and 2 plot the discretisation and normalisation errors as a function of the number of qubits used to prepare the distributions for the various propagators. Considering the $\sqrt{s_{max}}=100\,\text{GeV}$ distribution first, with the aim of allowing for classical simulations later, while keeping systematic errors relatively suppressed (i.e., minimising the number of qubits in the state-preparation circuit), we selected $n=6$, based on Fig. 1, which corresponds to errors in the range $\sim(10^{-3}-10^{-4})$. For the $\sqrt{s_{max}}=200\,\text{GeV}$ distributions we chose $n=9$, such that errors are in the $\sim(10^{-4}-10^{-5})$ range.
@@ 3.1 Variational method
Briefly summarising the method, the loss function that is minimised in the training using a classical optimiser is the $L_{\alpha}$ norm (for $\alpha=1,2,\infty$) between the target state, |$\ket{\text{tar}}$, and the generated state, |$\ket{\text{gen}}$,
@@
|$\ket{\text{gen}}=\ket{\psi(\vec{\theta})}=U(\vec{\theta})\ket{0^{n}}.$ (27)
As is well understood from the literature, variational methods are plagued by issues of trainability, specifically the gradients of the cost function vanishing exponentially in the size of the system (number of qubits), known as barren plateaus. Barren plateaus have been shown to be directly related to the expressitivity of the circuit ansätze used in the training [56]. In general, there is therefore a balance between allowing the circuit to be expressive enough to contain the desired solution, whilst also limiting the effects of barren plateaus. Given this, it is thought that variational approaches for state preparation will likely not be scalable for large system sizes. However, for the BW distributions considered in this article—which are all small-scale circuits—variational methods should give good results in practice. Indeed, no particular effort is made here to limit the expressivity of the circuit ansatz (for example by using some alternative ansätze such as the ones discussed in Refs. [57, 58]), as it was not considered necessary for these small systems.
@@ 3.2 Fourier expansion method
To again briefly summarise the method, the target $f$, is approximated by mapping to the interval $[-1,1]$ [^11] Specifically, it is represented in the interval $[0,1]$ and then a periodic extension is applied to map onto the interval $[-1,0)$. with a finite Fourier series of the form
@@
The potential benefits of this method is that, in contrast to the previous method, it should scale well to larger system sizes. However, one important consideration is that the method relies on a considerable number of ancilla qubits, and is a probabilistic state preparation; this is because the LCU method requires $\text{log}_{2}(2d+1)$ additional ancilla qubits, and to then prepare the correct state, all of these ancilla qubits must be post selected to give the zero state, which has success probability $p_{\text{success}}=\frac{\sum_{x}\left|f_{d}(x)\right|^{2}}{2^{n}(\sum_{k}\left|c_{k}\right|)^{2}}$. In practice, $p_{\text{success}}$ can be increased using amplitude amplification, although this of course means additional quantum resources. It is also worth noting that, in practice, the chosen value of $d$ dictates the final accuracy of the prepared distribution (i.e., larger $d$ means a smaller state-preparation error and JSD).
@@ 3.3 Results and comparison
Both the methods discussed in the previous sections were used in order to produce circuits that prepared BW distributions for the W, Z, and t resonances based on the $\sqrt{s_{max}}$ ranges and corresponding numbers of qubits in the circuit ansatz. For comparing the different methods, the metrics that are considered are the resources required for the circuit, in terms of the total number of qubits, $n$, alongside the number of 1- and 2-qubit gates, $g_{1q}$ and $g_{2q}$, respectively, and the accuracy of the prepared distribution, in terms of the state-preparation error, $\epsilon_{s}^{\text{CMSE}}$, and the JSD. In this section we will explicitly discuss the $9$-qubit circuits corresponding to $\sqrt{s_{max}}=200\,\text{GeV}$, in order to perform a comparison of the two methods for all resonances. However, results for the other COM energy range are also briefly discussed later.
@@
The Fourier expansion method was then analysed. For each of the resonances we produced a circuit that was roughly equivalent to the best one produced by the variational method in terms of the accuracy of the prepared distribution. In addition, we also produced another circuit where the accuracy of the prepared distribution was significantly greater than in the variational case (i.e., with an increased value of $d$). This was done to demonstrate that the method enables the preparation of larger-scale circuits while maintaining accuracy.
Figure 3 gives the generated distributions for each resonance using the variational method, while Figs. 4 and 6 give the generated distributions using the Fourier method for each of the resonances, respectively. Table 1 lists the metrics for the distributions produced for all COM ranges for both methods. Figure 12 in Section A.1 also plots the distributions generated by the variational and Fourier expansion methods for the $n=6$, $\sqrt{s_{max}}=100\,\text{GeV}$ circuits.
@@
Figure 3: True (blue triangles) and generated (orange circles) points for 9-qubit circuits for the BW distribution for various resonances up to $\sqrt{s_{max}}=200\,\text{GeV}$, generated using the variational method.
@@
Figure 4: True (blue triangles) and generated (orange circles) points for the W-boson BW distribution up to $\sqrt{s_{max}}=200\,\text{GeV}$, generated using the Fourier expansion method, for two different 9-qubit circuits corresponding to (left) a less accurate generated distribution that matches the accuracy of the variational method, and (right) a more accurate generated distribution.
@@
Figure 5: True (blue triangles) and generated (orange circles) points for the Z-boson BW distribution up to $\sqrt{s_{max}}=200\,\text{GeV}$, generated using the Fourier expansion method, for two different 9-qubit circuits corresponding to (left) a less accurate generated distribution that matches the accuracy of the variational method, and (right) a more accurate generated distribution.
@@
Figure 6: True (blue triangles) and generated (orange circles) points for the t-quark BW distribution up to $\sqrt{s_{max}}=200\,\text{GeV}$, generated using the Fourier expansion method, for two different 9-qubit circuits corresponding to (left) a less accurate generated distribution and (right) a more accurate generated distribution.
Table 1: Comparison of the metrics for the prepared distributions for each resonance for both COM energies, using either variational or Fourier expansion methods. “Res” stands for “Resonance”. For “Accuracy”, “Optimised” refers to the variationally-trained circuits with optimised hyperparameters, and “Matched” and “More” refer to the Fourier-expansion-method circuits, with either similar, or much greater accuracy, to the equivalent variationally trained circuits, respectively.
@@
[…]
Z Optimised $9$ $234$ $200$ $4.77\times 10^{-3}$ $0.024$ N/A
t Optimised $9$ $126$ $104$ $8.16\times 10^{-3}$ $0.070$ N/A
@@
It is worth also discussing that, for the $n=6$ circuits (with distributions plotted in Fig. 12 in Section A.1), we find that the variational method produces slightly more accurate results than the Fourier expansion method, and again requires significantly less resources. This indeed makes sense, given that we expect variational methods to perform well for small-scale circuits; however, one would expect the performance to decrease as the system size increases, which is indeed what we see when considering the $n=9$ circuits discussed previously.
As an aside, if we were to consider an example of a potential more realistic use case corresponding to an energy scale comparable to a modern collider experiment—where in reality it would be essential to really suppress systematic errors—of $\sqrt{s_{max}}=1\,\text{TeV}$, for example, then we can see from Fig. 7 that we would need around $n=16$ for systematic errors to be suppressed to around $\sim(10^{-5}-10^{-6})$. For circuits this size, variational methods are likely not to be performant due to the trainability issues previously discussed. On the contrary, the Fourier expansion method should still be effective for generating accurate distributions at this circuit size, although the resources required would likely be considerable.
As discussed previously, state-preparation for probability distributions is an active research topic, and future developments should hopefully allow for the preparation of accurate BW distributions with reduced circuit sizes, paving the way for less resource-intensive QMCI for HEP. We now move on to discuss an example application of the QMCI engine for calculating a cross-section in HEP, making use of the $6$-qubit, $\sqrt{s_{max}}=100\,\text{GeV}$, variationally-trained circuit for the W-boson BW distribution discussed in this section.

### v2

@@ 2.4 Implementing a generic cross-section calculation
In the introduction, we have mentioned so called generic building blocks for the computation of cross sections. To illustrate this, we provide here two concrete examples. From Eqs.˜1 and 4, a cross section with only a single integration variable reduces to
@@
where $s$ is the maximal value of $x$. In order to cast the above equation into the more rigorous one of the MCI problem as in Eq.˜6, noting that we are free to choose how to split the integrand for the expectation value into the product of function applied, $g(x)$, and probability distribution, $f_{X}(x)$, it follows that we can choose the function applied to be
@@
(i.e. the numerator). The probability distribution to sample from can then be associated to the propagator function (i.e. the denominator), such that
@@
The reason we chose this particular factorisation is twofold. First, as mentioned previously, implementing moments or products of moments of random variables as the function applied is straightforward to implement based on the QMCI engine’s existing functionality, and the RMSE scaling of the corresponding Fourier series’ are likely favourable as compared to those of more complex functions. Second, because the probability distribution corresponds to a real distribution that has physical significance (being used to model resonance behaviour in general), we can envision that future development of, for example, specific state-preparation methods tailored to this distribution [^12] Note that in this article we only consider general state-preparation methods for preparing the BW distribution, and we defer research into more tailored methods for future work. would not only be useful for this particular application, but also for future research directions in HEP (e.g. for quantum-computing applications where resonance behaviours are required to be modelled).
We can identify the general form of Eq.˜14 as an (unnormalised) relativistic BW distribution [47] (hereafter simply denoted by BW distribution). As mentioned previously, state preparation of arbitrary probability distributions is a difficult problem. However, in this case we need only prepare one specific distribution, and as will be discussed in the following section, it will suffice to prepare this distribution a single time (for a given COM energy) for each of the (small number of) relevant massive particles in the Standard Model, i.e., the W and Z bosons, the top (t) quark, and the Higgs boson. In the present example, the two building blocks are the function, $g$, which is a monomial, and the BW distribution, $f_{X}$.
@@
Such terms would then be absorbed in the monomials in the numerator, possibly with negative exponents. For this case of negative exponents, because the functions applied for FQMCI must obey certain smoothness conditions, as detailed previously in Section˜2.3.1, then in this case one would not be able to exploit FQMCI. Instead, an approach similar to those of prior works would be required, where the entire integrand is prepared as the probability distribution and then QAE run directly on this state (equivalent in our language to the ‘function applied’ being constant).
@@
where the term $x^{n}y^{m}$ means that the numerator can be made of any monomials which are a product of $x$ and $y$ raised to various powers. It implies that in order to match the form of Eq.˜8, the separable functions applied should be
@@
The QMCI engine provides efficient Fourier decomposition circuits for moments or products of moment of random variables, as discussed previously, and allows the user to provide the probability distribution to be sampled from as an input. In addition, when constraints on the phase space are required to be imposed on the integration, as given by the cut function, $C$, in Eq.˜5, then for simple constraints such as sums, maximum/minimum etc. the engine’s built-in thresholding operations can be used to achieve this. Thus—providing there are efficient ways to prepare BW distributions as quantum states, something discussed in the next section—in principle the QMCI engine can be used to perform quantum integration of a generic cross section in HEP.
It is worth noting that the cut function, $C$, is a function of the integration variables, and as discussed previously, possibly non-analytical. This means that in order to implement the restriction of phase space, the value $C(\Phi)$ should be numerically evaluated at each phase-space point. Such an approach has been followed in Section˜4 to set the boundaries of integration of a given variable that depends on another variable of integration; however, in this case the expression was analytical, and the quantum arithmetic implemented via the QMCI engine’s enhanced $P$-builder. For the case of some non-analytical form, one possibility is to implement the required quantum arithmetic by constructing the equivalent classical circuit, and then making it a reversible circuit. However, we note that this approach is unlikely to be efficient, and indeed, in the future there may be better ways of constructing such circuits. We leave this aspect for future work.
3 State preparation of relativistic BW distributions
We see that to be able to perform generic cross-section calculations using the QMCI engine, we require a methodology for preparing on a quantum register BW distributions for the massive propagators W, Z and t, [^13] In practice, completeness would also require a propagator for the Higgs boson. However, due to the narrow width of the resonance, generating this peak with sufficient resolution is highly challenging, given current qubit counts and methods. At high-energy colliders, all leptons are typically considered as massless, alongside all quarks (apart from the top quark), as their masses are negligible at these energy scales. In certain cases, the mass of the bottom quark is also included in theoretical calculations, but with the width assumed to be zero. as these constitute one of the building blocks of the method. As discussed briefly in the introduction, state preparation of probability distributions for QMCI is an active research topic, and several methods have been proposed in the literature (see, e.g. Refs. [48, 49, 50, 51, 52, 53, 54]); however, there is currently no standard methodology that is commonly used.
@@ 3 State preparation of relativistic BW distributions
[…]
@@
  • State-preparation error: this error is distinct from the other errors discussed as it does not arise from the limited number of qubits, but rather the imperfect method of preparing the distribution as a quantum state i.e. there is a difference between the ideal quantum state and the approximated state that is actually prepared. The state-preparation error can be defined in a variety of different ways. In this article, we consider the cumulative-distribution-function (CDF) mean-squared error (CMSE) between the probabilities of the prepared state, $f^{p}_{X}(x_{i})$, and the true probabilities, $\tilde{f}_{X}(x_{i})$, as the metric representing the state-preparation error
@@
All of these systematic errors can be considered relevant for the cross-section calculations considered later in Section˜4.
Considering the discretisation and normalisation errors in particular, for a given number of qubits, $n$, there will be a sub-range of the full potential support of the distribution where these systematic errors are sufficiently minimised (below a chosen bound) to have a negligible impact on a QMCI calculation. When considering cross-section calculations for a given experiment, the integration is only ever performed from $s_{min}=0\,\text{GeV}^{2}$ to some $s_{max}=S\,\text{GeV}^{2}$. Therefore, in practice, it will not be necessary to prepare the BW distribution to span the full potential support range $s_{min}=0\,\text{GeV}^{2}$ to $s_{max}=+\infty\,\text{GeV}^{2}$]—which would require an infinite number of qubits in order to sufficiently suppress these errors. Thus, whilst qubit numbers are still limited, for practical applications it will instead suffice to generate a range of circuits that prepare BW distributions with supports spanning a range of different COM energies squared, with each prepared using a sufficient number of qubits as to sufficiently suppress these systematic errors.
It is also worth mentioning that while hadron colliders collide particles at energies of up to tens of TeV ($13.6\,\text{TeV}$ for the LHC), due to the suppression of the parton distribution functions at high energies, the accessible energy range is effectively reduced to a few hundreds of  GeV.
To this end, and for illustration purposes, in this article we consider preparing BW distributions for the W, Z, and t resonances, corresponding to a COM energy, $\sqrt{s_{max}}=200\,\text{GeV}$, and for just the W to $\sqrt{s_{max}}=100\,\text{GeV}$. The value, $\sqrt{s_{max}}=100\,\text{GeV}$, was chosen such that the circuits for the cross-section calculations in Sections˜4.2 and 4.3 (which only use the W-boson propagator) were sufficiently small to be classically simulable using state-vector methods, while $\sqrt{s_{max}}=200\,\text{GeV}$ was chosen in order to be able to generate distributions covering all of the resonances.
Figures˜1 and 2 plot the discretisation and normalisation errors as a function of the number of qubits used to prepare the distributions for the various propagators. Considering the $\sqrt{s_{max}}=100\,\text{GeV}$ distribution first, with the aim of allowing for classical simulations later, while keeping systematic errors relatively suppressed (i.e. minimising the number of qubits in the state-preparation circuit), we selected $n=6$, based on Fig.˜1, which corresponds to errors in the range $\sim(10^{-3}-10^{-4})$. For the $\sqrt{s_{max}}=200\,\text{GeV}$ distributions we chose $n=9$, such that errors are in the $\sim(10^{-4}-10^{-5})$ range.
@@ 3.1 Variational method
[…]
@@
[…]
As is well understood from the literature (see, e.g. Refs. [26, 27, 28, 29, 30, 31]), variational methods are plagued by issues of trainability, specifically the gradients of the cost function vanishing exponentially in the size of the system (number of qubits), known as barren plateaus. Barren plateaus have been shown to be directly related to the expressitivity of the circuit ansätze used in the training [57]. In general, there is therefore a balance between allowing the circuit to be expressive enough to contain the desired solution, whilst also limiting the effects of barren plateaus. Given this, it is thought that variational approaches for state preparation will likely not be scalable for large system sizes. However, for the BW distributions considered in this article—which are all small-scale circuits—variational methods should give good results in practice. Indeed, no particular effort is made here to limit the expressivity of the circuit ansatz (for example by using some alternative ansätze such as the ones discussed in Refs. [58, 59]), as it was not considered necessary for these small systems.
@@ 3.2 Fourier expansion method
To again briefly summarise the method, the target function,$f$, is approximated by mapping to the interval [1,1][-1,1] [^15] Specifically, it is represented in the interval $[0,1]$ and then a periodic extension is applied to map onto the interval [1,0)[-1,0). with a finite Fourier series of the form
@@
The potential benefits of this method is that, in contrast to the previous method, it should scale well to larger system sizes. However, one important consideration is that the method relies on a considerable number of ancilla qubits, and is a probabilistic state preparation; this is because the LCU method requires $\text{log}_{2}(2d+1)$ additional ancilla qubits, and to then prepare the correct state, all of these ancilla qubits must be post selected to give the zero state, which has success probability $p_{\text{success}}=\frac{\sum_{x}\left|f_{d}(x)\right|^{2}}{2^{n}(\sum_{k}\left|c_{k}\right|)^{2}}$. In practice, $p_{\text{success}}$ can be increased using amplitude amplification, although this of course means additional quantum resources. It is also worth noting that, in practice, the chosen value of $d$ dictates the final accuracy of the prepared distribution (i.e. larger $d$ means a smaller state-preparation error and JSD).
@@ 3.3 Results and comparison
Both the methods discussed in the previous sections were used in order to produce circuits that prepared BW distributions for the W, Z, and t resonances based on the $\sqrt{s_{max}}$ ranges and corresponding numbers of qubits in the circuit ansatz. For comparing the different methods, the metrics that are considered are the resources required for the circuit, in terms of the total number of qubits, $n$, alongside the number of one- and two-qubit gates, $g_{1q}$ and $g_{2q}$, respectively, and the accuracy of the prepared distribution, in terms of the state-preparation error, $\epsilon_{s}^{\text{CMSE}}$, and the JSD. In this section we will explicitly discuss the nine-qubit circuits corresponding to $\sqrt{s_{max}}=200\,\text{GeV}$, in order to perform a comparison of the two methods for all resonances. However, results for the other COM energy range are also briefly discussed later.
@@
The Fourier expansion method was then analysed. For each of the resonances we produced a circuit that was roughly equivalent to the best one produced by the variational method in terms of the accuracy of the prepared distribution. In addition, we also produced another circuit where the accuracy of the prepared distribution was significantly greater than in the variational case (i.e. with an increased value of $d$). This was done to demonstrate that the method enables the preparation of larger-scale circuits while maintaining accuracy.
Figure˜3 gives the generated distributions for each resonance using the variational method, while Figs.˜4, 5 and 6 give the generated distributions using the Fourier method for each of the resonances, respectively. Table˜1 lists the metrics for the distributions produced for all COM ranges for both methods. Figure˜12 in Section˜A.1 also plots the distributions generated by the variational and Fourier expansion methods for the $n=6$, $\sqrt{s_{max}}=100\,\text{GeV}$ circuits.
@@
Figure 3: True (blue triangles) and generated (orange circles) points for nine-qubit circuits for the BW distribution for various resonances up to $\sqrt{s_{max}}=200\,\text{GeV}$, generated using the variational method.
@@
Figure 4: True (blue triangles) and generated (orange circles) points for the W-boson BW distribution up to $\sqrt{s_{max}}=200\,\text{GeV}$, generated using the Fourier expansion method, for two different nine-qubit circuits corresponding to (left) a less accurate generated distribution that matches the accuracy of the variational method, and (right) a more accurate generated distribution.
@@
Figure 5: True (blue triangles) and generated (orange circles) points for the Z-boson BW distribution up to $\sqrt{s_{max}}=200\,\text{GeV}$, generated using the Fourier expansion method, for two different nine-qubit circuits corresponding to (left) a less accurate generated distribution that matches the accuracy of the variational method, and (right) a more accurate generated distribution.
@@
Figure 6: True (blue triangles) and generated (orange circles) points for the t-quark BW distribution up to $\sqrt{s_{max}}=200\,\text{GeV}$, generated using the Fourier expansion method, for two different nine-qubit circuits corresponding to (left) a less accurate generated distribution and (right) a more accurate generated distribution.
Table 1: Comparison of the metrics for the prepared distributions for each resonance for both COM energies, using either variational or Fourier expansion methods. ‘Res’ stands for ‘Resonance’. For ‘Accuracy’, ‘Optimised’ refers to the variationally-trained circuits with optimised hyperparameters, and ‘Matched’ and ‘More’ refer to the Fourier-expansion-method circuits, with either similar, or much greater accuracy, to the equivalent variationally trained circuits, respectively.
@@
$200\,\text{GeV}$ Variational W Optimised 99 $180$ $152$ $4.01\times 10^{-4}$ $0.029$ N/A
Z Optimised 99 $234$ $200$ $4.77\times 10^{-3}$ $0.024$ N/A
t Optimised 99 $126$ $104$ $8.16\times 10^{-3}$ $0.070$ N/A
@@
It is worth also discussing that, for the $n=6$ circuits (with distributions plotted in Fig.˜12 in Section˜A.1), we find that the variational method produces slightly more accurate results than the Fourier expansion method, and again requires significantly less resources. This indeed makes sense, given that we expect variational methods to perform well for small-scale circuits; however, one would expect the performance to decrease as the system size increases, which is indeed what we see when considering the $n=9$ circuits discussed previously.
As an aside, if we were to consider an example of a potential more realistic use case corresponding to an energy scale comparable to a modern collider experiment—where in reality it would be essential to really suppress systematic errors—of $\sqrt{s_{max}}=1\,\text{TeV}$, for example, then we can see from Fig.˜7 that we would need around $n=16$ for systematic errors to be suppressed to around $\sim(10^{-5}-10^{-6})$. For circuits this size, variational methods are likely not to be performant due to the trainability issues previously discussed. On the contrary, the Fourier expansion method should still be effective for generating accurate distributions at this circuit size, although the resources required would likely be considerable.
As discussed previously, state-preparation for probability distributions is an active research topic, and future developments should hopefully allow for the preparation of accurate BW distributions with reduced circuit sizes, paving the way for less resource-intensive QMCI for HEP. We now move on to discuss an example application of the QMCI engine for calculating a cross-section in HEP, making use of the six-qubit, $\sqrt{s_{max}}=100\,\text{GeV}$, variationally-trained circuit for the W-boson BW distribution discussed in this section.